OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation 4x - 3y = 24. x intercept = (6, 0); y intercept = (0, -8) x intercept = (-6, 0); y intercept = (0, 8) x intercept = (-8, 0); y intercept = (0, -6) x intercept = (8, 0); y intercept = (0, 6)

Parth (parthkohli):

when \(x =0\), then the solution for \(y\) is the \(y\)-intercept. when \(y = 0\), then the solution for \(x\) is the \(x\)-intercept.

OpenStudy (anonymous):

just set y=0 and solve for x. then set x=0 and solve for y

Parth (parthkohli):

can you substitute \(y=0\) to get the \(x\)-intercept?

OpenStudy (johnweldon1993):

turn from standard form into slope intercept form 4x - 3y = 24 -3y = 24 - 4x y = 24/-3 + 4/3x y = -8 + 4/3x now plug in 0 for x....and 0 for y to get each so 0 = 4/3x - 8 8 = 4/3x 24 = 4x x = 6 thats the x intercept and y = 4/3(0) - 8 y = -8 so x intercept = 6,0 and y intercept = -8,0